OpenStudy (anonymous):

For the function f(x) = (3 – 4x)², find f^–1. Determine whether f^–1 is a function.

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

oh, okay

hero (hero):

\[ f(x) = (3 - 4x)^2 \\y = (3 - 4x)^2 \\x = (3 - 4y)^2 \\\pm \sqrt{x} = 3 - 4y \\\pm \sqrt{x} - 3 = -4y \\\frac{\mp \sqrt{x} + 3}{4} = y \\y = \frac{\mp \sqrt{x} + 3}{4} \\f^{-1}(x) = \frac{\mp \sqrt{x} - 3}{4} \]

hero (hero):

Ugh

hero (hero):

I hate \(\LaTeX\)

OpenStudy (anonymous):

it's fine don't worry about it :)

hero (hero):

Everything is correct except the last equation should also be +3

OpenStudy (anonymous):

okey dokey! thanks